OpenStudy (anonymous):

Carroll High School has a total of 94 boys and girls who play sports. If the number of boys is 9 more than four times the number of girls, how many boys play sports at this high school? Answer 23 77 92 17

OpenStudy (anonymous):

9x=y, x+y = 94

OpenStudy (anonymous):

i do not think that is right, it says is 9 more meaning +9 more than 4* the number of girls.

OpenStudy (sasogeek):

y=9+4x x+y=94

OpenStudy (anonymous):

there you go decimal answer but still

OpenStudy (anonymous):

sry lol my bad its 17

OpenStudy (anonymous):

pressed the wrong thing :P

OpenStudy (sasogeek):

andres most of ur questions are simultaneous equations, ever took the pain to learn how to solve them?

OpenStudy (anonymous):

btw, without even doing calculations u can logically solve this problem. we know its for times more and a total of 94 students. 92 and 77 are out righ away. considering its more then 4 times more, 23 is out too as it goes over 94. so only right answer is 17. another trick if you do not understand or like the mathematial way :)